# On the MST-ratio: Theoretical Bounds and Complexity of Finding the Maximum

### v2

@@ 3 Maximum MST-ratio of Geometric Point Sets
We show that MAX-EMST-ratio problem is NP-hard. A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$. We rely on a result by Dekster and Wilker [11], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
Lemma 3.1.
[11, Theorem 2] For every positive integer $n>0$ there exists a real number $\lambda_{n}=\sqrt{1-\frac{1}{\Theta(n)}}$ such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\cdot\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
@@
Lemma 3.2.
For every weighted complete graph $G$ on $n$ vertices, there exists a positive integer $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
Proof 3.3.
Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$. Pick $N$ such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number in Lemma 3.1. Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$. $N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$. So $G^{\prime}$ satisfies the conditions of Lemma 3.1, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$.
Theorem 3.4.
The problem MAX-EMST-ratio is NP-hard.
Proof 3.5.
By Lemma 3.2, the instances in Theorem 2.1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount. Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio. This is because the total number of edges in the MSTs of the two colors is always two fewer than the total number of vertices. Moreover, the amount we add is $O(n)$, according to Lemmas 3.1,3.2. Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem in polynomial time.
@@
Conjecture 3.6.
The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST-ratio are NP-hard.
@@
Lemma 3.7 ([18]).
Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily pairwise distinct) paths in $T$ such that each edge of $T$ belongs to precisely two of these paths.
Theorem 3.8.
@@
To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ and $B_{Q}$ on the set of red vertices and the set of blue vertices of $Q$ respectively, such that $|R_{Q}\cup B_{Q}|$ is at most $(3-\frac{4}{n-1})|T|$. The way we construct $R_{Q}$ and $B_{Q}$ is as follows (see Figure 3).
  • We add every edge $e$ of $T$ that has both endpoints in red (in blue) to $R_{Q}$ (to $B_{Q}$).
@@
If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$. Assume that the diameter of $T$ is at least $4$. Let $T^{\prime}$ be the tree obtained from $T$ by removing all the leaves of $T$ and let $\ell$ and $\ell^{\prime}$ be the number of leaves of $T$ and $T^{\prime}$, respectively. Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$. Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$. Using Lemma 3.7, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths. Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
Therefore $W_{P^{*}}+W_{L}\geq\frac{4(|T|-W_{L})}{n-1}+W_{L}\geq\frac{4}{n-1}|T|.$
@@
Note that the above bound is the best we can achieve in metric spaces. In fact, for every odd $n$, there exist metric spaces with $n$ points (elements) for which the maximum MST-ratio is arbitrarily close to $3-\frac{4}{n-1}$. To see this, let $n=2k+1$, and define the point set $S=\{v_{1},\ldots,v_{n}\}$. For every $i<n$, we set the weight of $v_{n}v_{i}$ to be $1$. For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be $\varepsilon$ for a sufficiently small $\varepsilon>0$. Any other edge has weight $2$.
@@
We first argue that the weights are metric. The edges of weight $1$ are the edges incident to $v_{n}$, and the edges of weight $\varepsilon$ form a matching. Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight $1$ and has at most one edge of weight $\varepsilon$. Thus $(a,b,c)$ is either $(\varepsilon,1,1)$, or $(\varepsilon,2,2)$, or $(1,1,2)$, or $(2,2,2)$, and in all cases we have $c\leq a+b$. Observe that $|MST(S)|=\varepsilon\times k+1\times k=k(1+\varepsilon)$. If we color $v_{1},v_{3},\ldots,v_{n}$ by red and $v_{2},v_{4},\ldots,v_{n-1}$ by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$. Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k(1+\varepsilon)}=(3-\frac{4}{n-1})\frac{1}{1+\varepsilon}$.
Theorem 3.10.
@@
There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
Proof 3.11.
@@
By Theorem 3.8, the maximum EMST-ratio for an $n$-element point set is upper bounded by $3-\frac{4}{n-1}$. On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing a minimum spanning tree of the point set in ${O}(n^{2})$ (see [10, 19]), and then coloring the vertices of the two components of it after removing the shortest edge by red and blue. Finally, as $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$, this is an ${O}(n^{2})$ time $3$-approximation algorithm.
@@
We remark that a slightly improved running time for computing an EMST is provided in [3]. Moreover, one could achieve a weaker approximation factor $3.47$ in a simpler way using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577}<3.47$ on the MAX-EMST-ratio. Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e. $2$-MAX-EMST-ratio), the approximation factor in Theorem 3.10 can be improved to $2.427$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [9]), and for $n>12$ a coloring with EMST-ratio greater than one can be computed in polynomial time (see [13]). As can be observed, the coloring used to present the $3$-approximation is fairly straightforward. However, there seems to be room for improvement by identifying more mingled colorings.
@@
We introduce Bipartite Colorings as good candidates to approximate the maximum EMST-ratio of a point set. By a bipartite coloring of a point set, we mean considering one of the Euclidean minimum spanning trees and then partitioning the point set into two subsets labeled by colors in which every edge of the taken EMST is colorful. Notice that since the EMST of a point set is not necessarily unique, it may have more than one bipartite colorings, namely one for each EMST. Intuitively, bipartite colorings give well-distributed colorings of the point set that can be computed in ${O}(n^{2})$ time ($O(n\log n)$ in the plane, see [5]). Our experiments suggest that bipartite colorings give a very good approximation of the maximum EMST-ratio for random point sets in the plane (see Section 5).
@@
Although bipartite colorings seem to be effective, they do not always produce the maximum EMST-ratio. Here is an extreme example: Consider $n=2k+1$ points on a zigzag path in the triangular grid and call them a Triangular Chain. Figure 4 illustrates a triangular chain for $k=5$. The maximum EMST-ratio of a triangular chain with $12k+$ 1 points is $\frac{3k-2}{2k}$ while the ratio for a bipartite coloring is $\frac{2k-1}{2k}$. (By stretching the chain slightly, one can ensure a unique EMST and, consequently, a unique bipartite coloring.) As $k$ increases, the proportion between the maximum EMST-ratio and this bipartite EMST-ratio approaches $1.5$.
@@
We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from a bipartite coloring for a point set in the plane. Recall that if this conjecture is correct, it would lead to an $O(n\log n)$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
Conjecture 3.12.
@@
Apart from the bipartite colorings, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple mixed coloring of a point set.
Question 3.13.
@@
The Average EMST-ratio of a point set $P$ is the average of EMST-ratio over all non-trivial $2$-colorings of $P$. It is automatically a lower bound for the maximum EMST-ratio. Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets. In fact, it is proved that the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit [12]. We present the proof for the generalized version of this result, i.e. for points within a $d$-dimensional unit cube (or any bounded region in ${{\mathbb{R}}}^{d}$).
Theorem 4.1.
@@
$|EMST(R)|\sim\beta(d).k^{1-1/d},\hskip 7.96677pt|EMST(B)|\sim\beta(d).(n-k)^{1-1/d},\hskip 7.96677pt|EMST(P)|\sim\beta(d).n^{1-1/d}$
with probability one, where $\beta(d)$ is the constant in (1), and $k_{0}$ is an appropriate constant. Since $n$ is significantly greater than $k_{0}$, and the EMST-ratio is at most $3$, the contribution of the EMST-ratios for colorings with $k<k_{0}$ red points or with $n-k<k_{0}$ blue points are negligible when computing the average EMST-ratio. Thus, the average EMST-ratio over all proper bi-partitions of $P$ is
$\sum_{P=R\cup B}\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}$ $\sim\sum_{k=k_{0}}^{n-k_{0}}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
@@
with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [6, 17]), and therefore converges to $d]{2}$ as $n$ goes to infinity.
Figure 5: A point set consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin. The corresponding minimum spanning tree is depicted. As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
Certainly, Theorem 3.8 implies that $3$ is an upper bound for the average EMST-ratio of point sets in Euclidean space. How much can this upper bound be improved? There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$. For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A\cup B\cup C\cup\{0\}$, where $A$ (resp. $B$, $C$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp. $b$,$c$), for $\varepsilon$ being small enough, see Figure 5. In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random bi-partition, each of the sets $A$, $B$ and $C$ will contain both colors.
As for the lower bound, we show that the average EMST-ratio for $n$ points is always at least $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [13].
Theorem 4.3.
@@
Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in a Euclidean MST, $T$, of $P$. Assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$. Considering a coloring $P=R\cup B$ of $P$ into red and blue points, let $i$ be the smallest index such that the edge $e_{i}$ is colorful. $EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$. This implies $|EMST(R)|+|EMST(B)|\geq|EMST(P)|-w_{i}$.

### v3

@@ 3 Maximum MST-ratio of Geometric Point Sets
We show that MAX-EMST-ratio problem is NP-hard. We remark that, a weighted complete graph does not necessarily correspond to the distance graph of a point set in $\mathbb{R}^{d}$. A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$.
We rely on a result by Dekster and Wilker [12], which shows that any weighted complete graph, whose weights belong to a small range, is realizable in Euclidean space.
Lemma 1
[12, Theorem 2] For every positive integer $n>0$ there exists a real number $\lambda_{n}=\sqrt{1-\frac{1}{\Theta(n)}}$ such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\cdot\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
@@
Lemma 2
Let $G$ be a weighted complete graph $G$ on $n$ vertices. Let $w_{1}$ and $w_{2}$ be the minimum and maximum weights of the edges of $G$ and let $\lambda_{n}$ be as defined in Lemma 1. For every $N$ such that $N\geq\max\{|w_{1}|,\frac{\lambda_{n}w_{2}-w_{1}}{1-\lambda_{n}}\}$, by increasing the weight of all edges in $G$ by $N$, we obtain a realizable graph in ${{\mathbb{R}}}^{n-1}$.
Proof
Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights in which the minimum and maximum weights are $N+w_{1}$ and $N+w_{2}$. As $N>\frac{\lambda_{n}w_{2}-w_{1}}{1-\lambda_{n}}$ then $N+w_{1}\geq\lambda_{n}(N+w_{2})$ and hence by Lemma 1, $G^{\prime}$ is realizable in ${{\mathbb{R}}}^{n-1}$.
Theorem 3.1
The MAX-EMST-ratio problem is NP-hard.
Proof
By Lemma 2, the instances of the MAX-MST-ratio problem with weights $1$ and $n$ can be turned into a geometric graph by increasing the weights of all edges by a fixed amount. Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio. This is because the total number of edges in the MSTs of the two colors is always $n-2$. Moreover, the amount we add is $O(n^{2})$, according to Lemma 2. Thus, from any instance of the MAX-MST-ratio with weights $1$ and $n$, we can construct an instance of the MAX-EMST-ratio problem in polynomial time and hence by Theorem 2.1.
@@
Conjecture 1
The $k$-MAX-EMST-ratio problem for $k\in\{2,3\}$ is NP-hard.
@@
Lemma 3 ([19])
Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily distinct) paths in $T$ such that each edge of $T$ is in precisely two of these paths.
Theorem 3.2
@@
To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ and $B_{Q}$ on the set of red vertices and the set of blue vertices of $Q$ respectively, such that $|R_{Q}\cup B_{Q}|$ is at most $(3-\frac{4}{n-1})|T|$. We construct $R_{Q}$ and $B_{Q}$ in the following way (see Figure 2).
  • We add every edge $e$ of $T$ with both endpoints in red (blue) to $R_{Q}$ ($B_{Q}$).
@@
If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$. Assume that the diameter of $T$ is at least $4$. Let $T^{\prime}$ be the tree obtained from $T$ by removing all the leaves of $T$ and let $\ell$ and $\ell^{\prime}$ be the number of leaves of $T$ and $T^{\prime}$, respectively. Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$. Furthermore, as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus, $\ell^{\prime}\leq\frac{n-1}{2}$. Using Lemma 3, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths. Hence, $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$. Therefore, it can be concluded that $W_{P^{*}}+W_{L}\geq\frac{4(|T|-W_{L})}{n-1}+W_{L}\geq\frac{4}{n-1}|T|.$
@@
Note that the above bound is the best we can achieve in metric spaces. In fact, for every odd $n$, there exist metric spaces with $n$ points (elements) for which the maximum MST-ratio is arbitrarily close to $3-\frac{4}{n-1}$. To see this, let $n=2k+1$, and define the point set $S=\{v_{1},\ldots,v_{n}\}$. For every $i<n$, we set the weight of $v_{n}v_{i}$ to $1$. For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be $\varepsilon$ for a sufficiently small $\varepsilon>0$. Any other edge has weight $2$.
@@
We first argue that the weights are metric. The edges of weight $1$ are the edges incident to $v_{n}$, and the edges of weight $\varepsilon$ form a matching. Therefore, any triangle with edge weights $a\leq b\leq c$ has zero or two edges of weight $1$ and has at most one edge of weight $\varepsilon$. Thus, $(a,b,c)$ is in $\{(\varepsilon,1,1),(\varepsilon,2,2),(1,1,2),(2,2,2)\}$, and in all cases we have $c\leq a+b$. Observe that $|\text{MST}(S)|=\varepsilon\times k+1\times k=k(1+\varepsilon)$. If we color $v_{1},v_{3},\ldots,v_{n}$ by red and $v_{2},v_{4},\ldots,v_{n-1}$ by blue, then $|\text{MST}(R)|=k$ and $|\text{MST}(B)|=2(k-1)=2k-2$. Thus, $|\text{MST}(R)|+|\text{MST}(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k(1+\varepsilon)}=(3-\frac{4}{n-1})\frac{1}{1+\varepsilon}$.
Theorem 3.3
@@
There exists an ${O}(n^{2})$-time $3$-approximation algorithm for the MAX-EMST-ratio problem.
Proof
@@
By Theorem 3.2, the maximum EMST-ratio for an $n$-element point set is upper bounded by $3-\frac{4}{n-1}$. On the other hand, one can achieve an EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing a minimum spanning tree of the point set in ${O}(n^{2})$ (see [11, 20]), and then coloring the vertices of the two components of it after removing the shortest edge by red and blue. Finally, as $\frac{n-2}{n-1}\geq\frac{(3-\frac{4}{n-1})}{3}$, this is an ${O}(n^{2})$ time $3$-approximation algorithm.
@@
We remark that a slightly improved running time for computing an EMST is provided in [3]. Moreover, one could achieve a weaker approximation factor $3.47$ in a simpler way using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577}<3.47$ on the MAX-EMST-ratio. Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e. $2$-MAX-EMST-ratio), the approximation factor in Theorem 3.3 can be improved to $2.427$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [10]), and for $n>12$ a coloring with EMST-ratio greater than $1$ can be computed in polynomial time (see [14]). As can be observed, the coloring used to present the $3$-approximation is fairly straightforward. However, there seems to be room for improvement by identifying more mingled colorings.
@@
We introduce Bipartite Colorings as good candidates to approximate the maximum EMST-ratio of a point set. By a bipartite coloring of a point set, we mean considering one of the Euclidean minimum spanning trees and then partitioning the point set into two subsets labeled by colors in which every edge of the taken EMST is colorful. Note that since the EMST of a point set is not necessarily unique, it may have more than one bipartite coloring, namely, one for each EMST. Intuitively, bipartite colorings give well-distributed colorings of the point set that can be computed in ${O}(n^{2})$ time ($O(n\log n)$ in the plane, see [6]). Our experiments suggest that bipartite colorings give a very good approximation of the maximum EMST-ratio for random point sets in the plane (see Section 0.A.3).
@@
Although bipartite colorings seem to be effective, they do not always produce the maximum EMST-ratio. Here is an extreme example: Consider $n=2k+1$ points on a zigzag path in the triangular grid and call them a Triangular Chain (see Figure 3). By stretching the chain slightly, one can ensure a unique EMST and hence a unique bipartite coloring. The maximum EMST-ratio of a triangular chain with $12k+$ 1 points is $\frac{3k-2}{2k}$ (The proof is provided in Section 0.A.1), while the ratio for a bipartite coloring is $\frac{2k-1}{2k}$ As $k$ increases, the proportion between the maximum EMST-ratio and this bipartite EMST-ratio approaches $1.5$.
@@
We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from a bipartite coloring for a point set in the plane. If this conjecture is correct, it would lead to an $O(n\log n)$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
Conjecture 2
@@
Apart from the bipartite colorings, we are interested in any improved approximation algorithm for the MAX-EMST-ratio that can be obtained by a simple mixed coloring of a point set.
Question A
@@
The Average EMST-ratio of a point set $P$ is the average of EMST-ratio over all proper $2$-colorings of $P$. It is automatically a lower bound for the maximum EMST-ratio. Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets. In fact, it is proved that the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit [13]. We present the proof for the generalized version of this result, i.e. for points within a $d$-dimensional unit cube (or any bounded region in ${{\mathbb{R}}}^{d}$).
Theorem 4.1
@@
$|\text{EMST}(R)|\sim$ $\beta(d).k^{1-1/d},\hskip 5.69054pt|\text{EMST}(B)|\sim\beta(d).(n-k)^{1-1/d},$
$|\text{EMST}(P)|\sim\beta(d).n^{1-1/d}$
with probability approaching $1$, where $\beta(d)$ is the constant in (1), and $k_{0}$ is an appropriate constant. Since $n$ is significantly greater than $k_{0}$, and the EMST-ratio is at most $3$, the contribution of the EMST-ratios for colorings with $k<k_{0}$ red points or with $n-k<k_{0}$ blue points are negligible when computing the average EMST-ratio. Thus, the average EMST-ratio over all proper bi-partitions of $P$ is
$\sum_{P=R\cup B}\frac{|\text{EMST}(R)|+|\text{EMST}(B)|}{|\text{EMST}(P)|}$ $\sim\sum_{k=k_{0}}^{n-k_{0}}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
@@
with probability $1$ as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [7, 18]), and therefore converges to $d]{2}$ as $n$ goes to infinity.
Refer to caption$(a)$Refer to caption$(b)$
Figure 4: (a) A point set whose average EMST-ratio approaches approximately $2.154$.
(b) A point set with average EMST-ratio $<1$. The point set consists of $n-5$ points placed in a small circle (the core) and $5$ outer points as vertices of a regular pentagon inscribed in the unit circle around the origin. The corresponding MST is depicted.
Certainly, Theorem 3.2 implies that $3$ is an upper bound for the average EMST-ratio of point sets in Euclidean space. How much can this upper bound be improved? There exist point sets in ${{\mathbb{R}}}^{2}$ with an average EMST-ratio $\approx 2.154$. For instance, consider the unit circle centered at the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A\cup B\cup C\cup\{0\}$, where $A$ (resp. $B$, $C$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp. $b$, $c$), for $\varepsilon$ being small enough (see Figure 4.a). In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random bi-partition, each of the sets $A$, $B$, and $C$ will contain both colors. Regarding the lower bound, we show that the average EMST-ratio for $n$ points is always at least $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [14].
Theorem 4.2
@@
Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges of an EMST of $P$, called $T$. Assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$. Considering a coloring $P=R\cup B$ of $P$ into red and blue points, let $i$ be the smallest index such that the edge $e_{i}$ is colorful. As $\text{EMST}(R)\cup\text{EMST}(B)\cup\{e_{i}\}$ is a spanning tree of $P$, it holds that $|\text{EMST}(R)|+|\text{EMST}(B)|\geq|\text{EMST}(P)|-w_{i}$.